### v2

@@ 5.3 SHACL extensions
The SHACL language consists of two main components: SHACL core and SHACL-SPARQL. Additionally, there exist non-standard SHACL extensions such as the DASH Data Shapes Vocabulary [47] and SHACL Advanced Features [48], all of which rely heavily on SPARQL. This paper focuses on SHACL core for several reasons. First, studies that extract constraints from RDF graphs primarily use SHACL core components. Notably, research by Spahiu et al. [28] and Rabbani et al. [29] indicates that SHACL can enhance DQA. Thus, we aim to determine if SHACL core alone is sufficient for this purpose. Second, while SHACL can be implemented in various programming languages (e.g., jena-shacl in Java and PySHACL in Python), each with their own optimizations, once SPARQL-based constraints are used, performance depends on the underlying query engine. Furthermore, while the SHACL recommendation details its extension via SPARQL, other languages could be used. Lastly, we choose to focus on SHACL core because it is typically easier to understand and use than writing complete SPARQL queries, allowing for more intuitive formulation of constraints accessible to domain experts [49]. Appendix D provides more details on SHACL extensions.
5.4 Strengths and capabilities of SHACL core
Despite the limitations to cover DQA in its entirety, SHACL core has many strengths for specific aspects of DQA. For example, it can perform syntactic validation, including pattern validation, datatype checks, and verifying whether values fall within specified ranges or lists. It can validate consistency between the data and vocabulary or ontology definitions, such as ensuring correct use of a property’s domain or range. It can also support best practices for publishing Linked Data, such as the use of hash URIs for entities, labels in resources (i.e., entities, classes, and properties), and some characteristics of URI designs (i.e., short URIs and no parameters). Moreover, SHACL can verify the correct application of property characteristics such as irreflexive, functional, asymmetric, and inverse-functional. Finally, it is well-suited for defining the expected structure of class instances, specifying required properties, their types, and other basic constraints [13].
@@ 6 Conclusion and Future Work
The DQ metrics defined by Zaveri et al. [1] may benefit from refinement, as some are overly specific (e.g., those that require spatial data representations) and difficult to generalize, limiting their practical use in KG DQA.
We conclude that assessing DQ is a much broader concept than merely checking constraints, as it encompasses the data source, processing system, downstream tasks, and human factors [50]. SHACL core focuses on data graphs, limiting its applicability for comprehensive DQA. Future work should explore a broader DQA framework that also assesses semantics and external artifacts. Additionally, DQ is inherently context-dependent [15], and as we’ve seen, many metrics require domain knowledge. Regardless of whether SHACL shapes or other mechanisms are used, a key direction is to find systematic ways to represent and extract the domain knowledge needed for these types of metrics.
@@ Appendix A Related Work
Table 6 presents a summary of DQA tools proposed in the literature, including their name, input type, the DQ dimensions or aspects they cover, and whether they are currently maintained.
| Tool | Input | Dimensions/Aspects | Maintained |
| KGHeartbeat [23] | Dump/SPARQL | Multiple dimensions from [1] | Yes |
@@
| Luzzu [21] | Dump/SPARQL | Multiple dimensions from [1] | No |
@@ B.1.1 Availability
According to Zaveri et al. [1] , the Availability dimension refers to the degree to which a dataset, or part of it, is accessible, retrievable, and ready for use. Five metrics are proposed for this dimension, most of which require accessing resources on the web (e.g., checking the dereferenceability of URIs or detecting dead or broken links). From the five metrics, only one could be partially checked, and it’s the one that refers to the availability of an RDF dump (A2), and whether it can be downloaded. For this metric, we defined shape SHLABEL:shape:availability_void_shape, which checks if there’s an RDF dump provided, by verifying the existence of the property void:dataDump in the VoID description of the dataset. This can also be checked for the property dcat:downloadURL from the DCAT vocabulary. However, in both cases, SHACL cannot verify if the referenced dump can be downloaded, as it cannot perform network requests. Therefore, we state that this metric can be partially covered with SHACL core.
@@ B.1.3 Interlinking
Interlinking describes the extent to which entities representing the same concept are connected, either within a single data source or across multiple sources, for which the Zaveri et al. [1] identified three metrics (I1-I3).
@@ B.2.1 Syntactic Validity
Each defined shape for the aspects SV2A1 and SV2A2 must be instantiated with a specific property. Moreover, each shape targets subjects of the specified property; hence, the validation report will flag a violation for each subject that uses these properties with a value that is not allowed (SHLABEL:shape:accuracy_allowed_values_range, SHLABEL:shape:accuracy_allowed_values_list, doesn’t follow the specified pattern SHLABEL:shape:accuracy_syntactic_rules), or lacks the specified value (SHLABEL:shape:accuracy_allowed_values_has_value). Note that for all these shapes, we need to consider assumption (A3), as we need domain knowledge to instantiate the shapes with the allowed values (SHLABEL:shape:accuracy_allowed_values_range, SHLABEL:shape:accuracy_allowed_values_list, SHLABEL:shape:accuracy_allowed_values_has_value) or pattern (SHLABEL:shape:accuracy_syntactic_rules).
The third aspect (SV2A3) verifies if values conform to the specific “RDF pattern” and that the “types” are defined for specific resources. Upon exchange with the authors of [1], we were able to capture the meaning of “RDF pattern” as “the expected structural and semantic patterns in the data that arise from community practice”. For example, while the VoID vocabulary does not define a specific property to specify the homepage of a dataset, it is common practice to use foaf:homepage for this purpose. This is something that arises from community practice, but it’s not stated in the formal definition of the vocabulary. In this case, SHACL core can cover this aspect, because shapes can specify the expected structure for class instances. We illustrate a shape for this aspect in SHLABEL:shape:accuracy_rdf_pattern; however, note that this depends on the vocabulary, and it’s restricted to stating the class and expected (or commonly used) properties. Moreover, for this shape, it is necessary to consider assumption (A3), since to instantiate this shape one needs knowledge about community practices.
@@
Overall, we argue that the SHACL core can partially cover the metric SV2, as we were unable to define shapes to check all aspects of this metric.
The last metric (SV3) identifies malformed literals, which can be literals that are members of another datatype or that don’t conform to the lexical syntax of their datatype. In this case, we defined SHLABEL:shape:accuracy_malformed_literal, which enforces the constraint sh:datatype, that specifies the expected datatype for the values of the property. This constraint also raises a violation if the value is an ill-typed literal, for the datatypes supported by SPARQL 1.1 [25].
@@ B.2.2 Semantic Accuracy
The second metric (SA2) verifies that there are no inaccurate values by checking three aspects: whether functional dependencies hold in the data (SA2A1), by comparing literal values of a resource (SA2A2), or via crowd-sourcing (SA2A3). The only way to compare values of different properties with SHACL core is by leveraging the Property Pair Constraint components. However, these constraints allow you to compare values between different properties when the triples share the same subject. To verify the aspect SA2A1, the comparison needs to be done across multiple entities; hence, triples won’t share the same subject. Additionally, the aspect SA2A3 can’t be covered with SHACL core, since SHACL is a declarative language, and crowd-sourcing involves human input. Therefore, for this metric, we were only able to define shape SHLABEL:shape:semantic_accuracy_inaccurate_values for the aspect SA2A2, which defines a constraint between literal values of a resource, using the constraint sh:equals as an example. Note that this shape can be modified to consider the other Property Pair constraints sh:lessThan, sh:lessThanOrEquals, or sh:disjoint.
@@ B.2.3 Consistency
Metric CN7 checks for no negative dependencies or correlation among properties by using association rules. This is defined in [56], where they first cluster the data into semantically correlated subsets and then they apply advanced techniques to identify correlations between properties, inverse relations, among others. The authors state that their tool is able to detect positive and negative association rules, which show occurrence dependencies among properties in a cluster. In particular, negative association rules state that properties on both sides of the association rule never appear together in instances of the cluster. An example of negative association rules is title , which specifies that no instance has both properties, name and title, at the same time. We consider that this metric can be partially addressed using SHACL core by translating association rules into expressions that use only logical operators, such as , , and , and then representing these expressions as SHACL constraints to validate whether the rules hold on the data. It is important to note that, while association rules are typically mined based on support and confidence from data, the translation presented here treats the rule as a strict logical constraint.
@@ B.2.4 Conciseness
Conciseness refers to minimizing redundancy at both the schema and data levels. For this dimension, Zaveri et al. [1] identified three metrics (CS1-CS3). The first metric (CS1) refers to high intensional conciseness, which can be measured by calculating the ratio (for the case of linked data, the schema is represented by the ontology or vocabularies used by the dataset) [51]. In this case, to identify non-unique properties and classes, we would need to recognize equivalent properties and classes with different names (e.g., a dataset can use the properties name and firstName. While they have different names, they refer to the same concept) [51]. However, this entails understanding the semantics of properties and classes to decide if they are equivalent in meaning, which is something SHACL core cannot do.
@@
In the first case, to obtain the number of unique objects, we would need to identify two instances that are equal but use different identifiers (in the case of linked data, this refers to the URI). To identify similar or equivalent entities, we need to check that there are no two entities with different URIs with the same (or similar) values for all their properties. However, this would require comparing property values across different entities, which SHACL core doesn’t support. Property Pair constraints only allow comparisons between properties of the same subject. Moreover, to determine the similarity between the values of properties, we would need to use some similarity measure, which is not something that can be done with SHACL core.
@@ B.3 Contextual
This section presents the shape definitions for the dimensions Relevancy, Trustworthiness, and Timeliness from the Contextual group. It also includes the remaining shape definitions for the dimension Timeliness, which were not included in Section 3.3.
@@ B.4.4 Interpretability
The validation report for SHLABEL:shape:interpretability_self_descriptive_formats outputs a violation for each entity that is not an IRI. In this case, the DQ measure is a ratio measure. For SHLABEL:shape:interpretability_self_descriptive_formats_properties, the validation report outputs a violation for every object of PROPERTY_URI that it’s not an IRI. In this case, the DQ measure is a composite measure.

### v3

@@ 5.3 SHACL extensions
The SHACL language consists of two main components: SHACL core and SHACL-SPARQL. Additionally, there exist non-standard SHACL extensions such as the DASH Data Shapes Vocabulary [48] and SHACL Advanced Features [49], all of which rely heavily on SPARQL. This paper focuses on SHACL core for several reasons. First, studies that extract constraints from RDF graphs primarily use SHACL core components. Notably, research by Spahiu et al. [28] and Rabbani et al. [29] indicate that SHACL can enhance DQ. Thus, we aim to determine if SHACL core alone is sufficient for this purpose. Second, while SHACL can be implemented in various programming languages (e.g., jena-shacl in Java and PySHACL in Python), each with their own optimizations, once SPARQL-based constraints are used, performance depends on the underlying query engine. Furthermore, while the SHACL recommendation details its extension via SPARQL, other languages could be used. Lastly, we choose to focus on SHACL core because it is typically easier to understand and use than writing complete SPARQL queries, allowing for more intuitive formulation of constraints accessible to domain experts [50]. Appendix D provides more details on SHACL extensions.
@@ 6 Conclusion and Future Work
The DQ metrics defined by Zaveri et al. [1] may benefit from refinement, as some are overly specific (e.g., those that require spatial data representations) and difficult to generalize, limiting their practical use in KG DQA. Moreover, new metrics (e.g., bias [51]) are not covered in the survey; we suggest extending this work to test whether SHACL core can cover them.
Assessing DQ goes beyond checking constraints, encompassing the data source, the system processing the data, downstream tasks, and human factors [52]. SHACL core focuses on data graphs, limiting its applicability for comprehensive DQA. Future work should explore hybrid approaches combining SHACL’s constraint checking with tools that access external resources and LLMs to overcome its limited semantic awareness. It would also be useful to identify which limitations are inherent to SHACL core design that can be addressed through extensions.
@@ Appendix A Related Work
Table 6 presents a summary of DQA tools proposed in the literature, including their name, input type, the DQ dimensions they cover, and whether they are currently maintained.
| Tool | Input | Dimensions | Maintained |
| KGHeartbeat [23] | Dump/SPARQL | Availability, Believability, Completeness, Conciseness, Consistency, Currency, Interlinking, Interoperability, Interpretability, Licensing, Performance, Representational Conciseness, Reputation, Security, Semantic Accuracy, Timeliness, Understandability, Verifiability, Versatility [1] | Yes |
@@
| Luzzu [21] | Dump/SPARQL | Representational Conciseness, Interoperability, Interpretability, Versatility, Understandability, Timeliness, Semantic Accuracy, Consistency, Conciseness, Availability, Licensing, Interlinking, Security, Performance [1] | No |
@@ B.1.1 Availability
According to Zaveri et al. [1] , the Availability dimension refers to the degree to which a dataset, or part of it, is accessible, retrievable, and ready for use. Five metrics are proposed for this dimension, most of which require accessing resources on the web (e.g., checking the dereferenceability of URIs or detecting dead or broken links). From the five metrics, only A2 can be partially checked, and it refers to the availability of an RDF dump, and whether it can be downloaded. For this metric, we defined shape SHLABEL:shape:availability_void_shape, which checks if there’s an RDF dump provided, by verifying the existence of the property void:dataDump in the VoID description of the dataset. This can also be checked for the property dcat:downloadURL from the DCAT vocabulary. However, in both cases, SHACL cannot verify if the referenced dump can be downloaded, as it cannot perform network requests. Therefore, we state that this metric can be partially covered with SHACL core.
@@ B.1.3 Interlinking
Interlinking describes the extent to which entities representing the same concept are connected, either within a single data source or across multiple sources, for which Zaveri et al. [1] identified three metrics (I1-I3).
@@ B.2.1 Syntactic Validity
Each defined shape for the aspects SV2A1 and SV2A2 must be instantiated with a specific property. Moreover, each shape targets subjects of the specified property; hence, the validation report will flag a violation for each subject that uses these properties with a value that is not allowed (SHLABEL:shape:accuracy_allowed_values_range, SHLABEL:shape:accuracy_allowed_values_list), doesn’t follow the specified pattern (SHLABEL:shape:accuracy_syntactic_rules), or lacks the specified value (SHLABEL:shape:accuracy_allowed_values_has_value). Note that for all these shapes, we need to consider assumption (A3), as we need domain knowledge to instantiate the shapes with the allowed values (SHLABEL:shape:accuracy_allowed_values_range, SHLABEL:shape:accuracy_allowed_values_list, SHLABEL:shape:accuracy_allowed_values_has_value) or pattern (SHLABEL:shape:accuracy_syntactic_rules).
The third aspect (SV2A3) verifies if values conform to the specific “RDF pattern” and that the “types” are defined for specific resources. Upon exchange with the authors of [1], we were able to capture the meaning of “RDF pattern” as “the expected structural and semantic patterns in the data that arise from community practice”. For example, while the VoID vocabulary does not define a specific property to specify the homepage of a dataset, it is common practice to use foaf:homepage for this purpose. This is something that arises from community practice, but it’s not stated in the formal definition of the vocabulary. In this case, SHACL core can cover this aspect, because shapes can specify the expected structure for class instances. We illustrate a shape for this aspect in SHLABEL:shape:accuracy_rdf_pattern; however, note that this depends on the vocabulary, and it’s restricted to stating the class and expected (or commonly used) properties. Moreover, for this shape, it is necessary to consider assumption (A3), since to instantiate this shape we need knowledge about community practices.
@@
Overall, we argue that SHACL core can partially cover the metric SV2, as we were unable to define shapes to check all aspects of this metric.
The last metric (SV3) identifies malformed literals, which can be literals that are members of another datatype or that don’t conform to the lexical syntax of their datatype. In this case, we defined SHLABEL:shape:accuracy_malformed_literal, which enforces the constraint sh:datatype, that specifies the expected datatype for the values of the property. This constraint also raises a violation if the value is an ill-typed literal, for the datatypes supported by SPARQL 1.1 [25]. For datatypes not supported by SPARQL 1.1, shape SHLABEL:shape:accuracy_malformed_literal should be modified to use sh:pattern to define the expected value format (datatype).
@@ B.2.2 Semantic Accuracy
The second metric (SA2) verifies that there are no inaccurate values by checking three aspects: whether functional dependencies hold in the data (SA2A1), by comparing literal values of a resource (SA2A2), or via crowd-sourcing (SA2A3). The only way to compare values of different properties with SHACL core is by leveraging Property Pair Constraint components. However, these constraints allow you to compare values between different properties when the triples share the same subject. To verify the aspect SA2A1, the comparison needs to be done across multiple entities; hence, triples won’t share the same subject. Additionally, the aspect SA2A3 can’t be covered with SHACL core, since SHACL is a declarative language, and crowd-sourcing involves human input. Therefore, for this metric, we were only able to define shape SHLABEL:shape:semantic_accuracy_inaccurate_values for the aspect SA2A2, which defines a constraint between literal values of a resource, using the constraint sh:equals as an example. Note that this shape can be modified to consider the other Property Pair constraints sh:lessThan, sh:lessThanOrEquals, or sh:disjoint.
@@ B.2.3 Consistency
Metric CN7 checks for no negative dependencies or correlation among properties by using association rules. This is defined in [58], where they first cluster the data into semantically correlated subsets and then they apply advanced techniques to identify correlations between properties, inverse relations, among others. The authors state that their tool is able to detect positive and negative association rules, which show occurrence dependencies among properties in a cluster. In particular, negative association rules state that properties on both sides of the association rule never appear together in instances of the cluster. An example of negative association rules is , which specifies that no instance has both properties, and , at the same time. We consider that this metric can be partially addressed using SHACL core by translating association rules into expressions that use only logical operators, such as , , and , and then representing these expressions as SHACL constraints to validate whether the rules hold on the data. It is important to note that, while association rules are typically mined based on support and confidence from data, the translation presented here treats the rule as a strict logical constraint.
@@ B.2.4 Conciseness
Conciseness refers to minimizing redundancy at both the schema and data levels. For this dimension, Zaveri et al. [1] identified three metrics (CS1-CS3). The first metric (CS1) refers to high intensional conciseness, which can be measured by calculating the ratio (for the case of linked data, the schema is represented by the ontology or vocabularies used by the dataset) [53]. In this case, to identify non-unique properties and classes, we would need to recognize equivalent properties and classes with different names or identifiers (e.g., a dataset can use the properties name and firstName. While they have different names, they refer to the same concept) [53]. However, this entails understanding the semantics of properties and classes to decide if they are equivalent in meaning, which is something SHACL core cannot do.
@@
In the first case, to obtain the number of unique objects, we need to verify that there are no two instances which are equal but use different identifiers (in linked data, this refers to the URI). We would need to check that there are no two entities with different URIs with the same (or similar) values for all their properties. However, this would require comparing property values across different entities, which SHACL core doesn’t support. Property Pair constraints only allow comparisons between properties of the same subject. Moreover, to determine the similarity between the values of properties, we would need to use some similarity measure, which is not something that can be done with SHACL core.
@@ B.3 Contextual
This section presents the shape definitions for the dimensions Relevancy, Trustworthiness, and Understandability from the Contextual group. It also includes the remaining shape definitions for the dimension Timeliness, which were not included in Section 3.3.
@@ B.4.4 Interpretability
The validation report for SHLABEL:shape:interpretability_self_descriptive_formats outputs a violation for each entity that is not identified with an IRI. In this case, the DQ measure is a ratio measure. For SHLABEL:shape:interpretability_self_descriptive_formats_properties, the validation report outputs a violation for every object of PROPERTY_URI that it’s not an IRI. In this case, the DQ measure is a composite measure.